### v3

# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

Abstract
There has been a long-standing and widespread belief that Bayesian Optimization (BO) with standard Gaussian process (GP), referred to as standard BO, is ineffective in high-dimensional optimization problems. While this belief sounds reasonable, strong empirical evidence is lacking. In this paper, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems for high-dimensional optimization. We found that, surprisingly, when using Matérn kernels and Upper Confidence Bound (UCB), standard BO consistently achieves top-tier performance, often outperforming other BO methods specifically designed for high-dimensional optimization. Contrary to the stereotype, we found that standard GP equipped with Matérn kernels can serve as a capable surrogate for learning high-dimensional functions. Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also is robust in accommodating various structures within target functions. Furthermore, with standard GP, achieving promising optimization performance is possible via maximum a posterior (MAP) estimation with diffuse priors or merely maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models. In parallel, we also investigated and analyzed alternative popular settings in running standard BO, which, however, often fail in high-dimensional optimization. This might link to the a few failure cases reported in literature. We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems. The code for our experiment is available at https://github.com/XZT008/Standard-GP-is-all-you-need-for-HDBO.
1 Introduction
Many applications require optimizing complex functions, allowing queries only for the function values, possibly with noises, yet without any gradient information. Bayesian Optimization (BO) [28, 23] is a popular approach for addressing such challenges. BO typically employs Gaussian process (GP) [33] as a probabilistic surrogate. It iteratively approximates the target function, integrates the posterior information to maximize an acquisition function for generating new inputs at which to query, then updates the GP model with new examples, and concurrently approach the optimum.
Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems [8, 25, 6, 24, 18, 32, 19]. It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 25]. This belief may partly stem from the intuition that, as a kernel method, GPs struggle to handle high-dimensional inputs for covariance modeling, rending them incapable of providing an expressive prior for high-dimensional functions.
Kernel GP Training Acquisition Function
Matérn ✓ L-BFGS/MAP/Diffuse prior ✓ UCB ✓
ADAM/MAP/Diffuse prior ✓
L-BFGS/MAP/Concentrated prior ✗ EI ✗
SE ✗ ADAM/MAP/Concentrated prior ✗
L-BFGS/MLE ✗ TS ✗
ADAM/MLE ✓
Table 1: Favorable and unfavorable settings of running standard BO for high-dimensional optimization. SE: square exponential; EI: expected improvement; TS: Thompson sampling; MLE: maximum likelihood estimation.
A continuous line of research is dedicated to developing novel high-dimensional BO methods. The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs. One class of methods assumes a decomposition structure within the functional space [16, 26, 10, 34]. In other words, the target function can be expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables. Another family of methods assumes that the inputs are intrinsically low-rank. These methods either project the original inputs into a low-dimensional space [32, 25, 18] or use sparse-inducing priors to trim down massive input variables [6]. Subsequently, GP surrogates are built with the reduced input dimensions.
While the aforementioned beliefs and concerns may appear intuitive and reasonable, there lacks strong empirical evidence confirming that standard BO is indeed incapable of high-dimensional optimization. To bridge this gap, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems. Surprisingly, we found that, with ARD Matérn kernels and UCB as the acquisition function, the performance of standard BO consistently ranks among the best and often surpasses existing methods by a substantial margin. Those methods are specifically designed for high-dimensional optimization. This leads us to call for a re-evaluation and in-depth study of the potential of standard BO for high-dimensional optimization. The major discoveries are summarized as follows.
  • Optimization performance. We investigated BO with standard GP regression across eleven benchmarks widely used in literature, including six synthetic and five real-world high-dimensional optimization tasks. The number of optimization variables varies from 30 to 388. In comparison with six state-of-the-art high-dimensional BO methods, standard BO with ARD Matérn kernel and UCB consistently achieves the best (or close to the best) optimization performance across all the tasks. In contrast, the competing methods demonstrate strong performance when the target function aligns with their structural assumption. However, in cases of mismatch or when the structure of the target function is agnostic, such as those in real-world problems, the performance of the competing methods often experiences a significant drop. This underscores that BO with standard GP not only can excel in high-dimensional optimization but also is robust in accommodating various structures within the target functions.
  • Surrogate Learning. We investigated the task of learning high-dimensional functions with standard GP regression. Across ten benchmark datasets, our findings reveal that the prediction of standard GP equipped with Matérn kernels aligns well with the ground-truth, mostly performing better than more complex GP models that induce low-rank structures. This observation is in line with the recent results in kernel operator learning [3], where inputs are discretized functions — typically high-dimensional (hundreds or thousands of dimensions). It confirms that standard GP can serve as a capable surrogate for high-dimensional functions.
  • Computational Efficiency. We discovered that, with standard GP, achieving promising optimization performance is possible by using MAP estimation with diffuse priors or just maximum likelihood estimation. The marginalization of the parameters (kernel parameters and noise variance) via Markov-Chain Monte Carlo (MCMC) provides minor benefits while significantly increasing the training cost. In contrast, due to the usage of more complex surrogate models, training for alternative high-dimensional BO methods is often more expensive. Some of these methods heavily rely on extensive MCMC sampling to prevent overfitting, making the training of the surrogate model a computational bottleneck.
  • Failure Modes with Other Settings. We found failure modes of standard BO with other popular settings, such as using vanilla Square-Exponential (SE) kernels, Expected-Improvement (EI) and Thompson sampling (TS) as acquisition functions, and MAP estimation with concentrated Gamma priors over length-scale parameters. We conducted a careful ablation study to reveal the failure of these choices in both optimization and function learning, providing analysis and discussions. In Table 1, we list the discovered favorable and unfavorable settings for high-dimensional optimization.
2 Standard Bayesian Optimization
Consider maximizing a $d$-dimensional black-box objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$. The form of $f(\cdot)$ is unknown. We can only query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, and no gradient information is accessible. We aim to find ${\bf x}^{\dagger}=\operatornamewithlimits{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$. To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to learn and predict $f(\cdot)$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction. This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate. The utility typically encapsulates an exploration-exploitation trade-off. By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum. Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy. BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model. Then the iterative procedure repeats until BO converges to the optimum or a predefined stopping criterion is met.
The standard BO adopts GP regression [33] for surrogate modeling. Specifically, a GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is often chosen as a Mercer kernel function. For example, a popular kernel used in BO is the Matérn-5/2 kernel with Automatic Relevance Determination (ARD),
[…]
where $\tau=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\eta}})({\bf x}-{\bf x}^{\prime})}$, $a>0$ is the amplitude, and $\bm{\eta}>{\bf 0}$ are the length-scale parameters. We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter. Alternatively, one can use a single length-scale parameter shared by all the input dimensions.
Suppose we have collected a set of training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$. Due to the GP prior, the function values at ${\bf X}$, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Next, a Gauss noise model is used to model the likelihood of ${\bf y}$, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance. One can marginalize ${\bf f}$ out to obtain
$p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}).$ (2)
To estimate the GP parameters (e.g., $a$, $\bm{\eta}$, and $\sigma^{2}$), one can place a prior over these parameters and use MCMC to compute posterior samples or perform MAP estimation. One can also simply maximize (2) to perform MLE. The predictive distribution (given the GP parameters) has a closed form: at any new input ${\bf x}^{*}$, the predictive distribution is conditional Gaussian,
$p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right),$ (3)
where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
In each iteration, given the current GP surrogate, we maximize an acquisition function to identify the next input location at which to query. One commonly used acquisition function is the upper confidence bound (UCB), defined as
$\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X},$ (4)
where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) [14], Thompson sampling (TS) [27], among others.
3 High Dimensional Bayesian Optimization
An enduring and widespread belief is that when the input dimension $d$ is high (e.g., a few hundreds), the standard BO is prone to failure. This belief might partly arise from an intuition that commonly used kernels, such as in (1), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function. As a result, there is a dedicated line of research aimed at developing novel high-dimensional BO methods. The key idea of these methods is to introduce strong structural assumptions in surrogate modeling to sidestep directly handling the high-dimensional inputs in kernels.
3.1 Structural Assumption in Functional Space
The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
$f({\bf x})=\sum\nolimits_{j=1}^{M}f_{j}({\bf x}^{j}),\;\;f_{j}\sim\mathcal{GP}\left(m_{j}({\bf x}^{j}),\kappa_{j}(\cdot,\cdot)\right),$ (5)
where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$. There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$. In [16], the input ${\bf x}$ is partitioned into non-overlapped groups. After every a few BO iterations, a new partition is selected from a set of random partitions. The selection is based on the model evidence. In [26], the variable groups can overlap, and are represented as the maximum cliques on a dependency graph. A Gibbs sampling method was developed to learn the structure of the dependency graph. Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning. However, the most recent work [34] points out that, learning the group structure through relatively small data can be misleading. A wrongly learned structure can cause queries to be stuck at local optima. The work proposes RDUCB that randomly decomposed ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
3.2 Structural Assumption in Input Space
Another class of methods assumes a low-rank structure in the input space.
Low-dimensional Embedding. Wang et al., [32] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al., [25] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially imposes further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  [18]. ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue.
Partial Variable Dependency. Another strategy is to trim down the input variables. The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [11]. While one can also use MAP estimation, the performance is often much worse.
4 Comprehensive Evaluation
We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic experiments.
4.1 Experimental Settings
We employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
$\text{Fun}(d,d^{\prime})$ Structure
Ackley(300, 150) Partial variable dependency
Rosenbrok(300, 100) Partial variable dependency
Nonoverlap additive decomposition
Hartmann6(300,6) Partial variable dependency
Rosenbrock(100, 100) Nonoverlap additive decomposition
Stybtang(200, 200)
Ackley(150, 150) None
Table 2: Synthetic benchmarks and the structures within the target function: $d$ is the input dimension, and $d^{\prime}$ is the number of effective variables to compute the target function.
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section A.1 of the Appendix. Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The ground-truth of the target function values is computed with the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 2.
Real-World Benchmarks. We employed the following real-world benchmark problems.
  • Mopta08 (124)  [13], a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints. We followed [6] to encode the constraints as a soft penalty, which is added into the objective.
  • SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
  • Rover (60), a rover trajectory planning problem from [31]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
  • DNA (180) [35], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [22]. Prior analysis [35] shows that only around 43 variables are relevant to the target function.
  • NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
Method Structural Assumption
ALEBO, HESBO Low-dim embedding
SaasBO Partial variable dependency
RDUCB, Tree UCB Additive function decomposition
SBO, TURBO, BNN None
Table 3: BO methods and their structural assumption about the target function.
Methods. To execute the standard BO (SBO), we used GP regression with ARD Matérn-5/2 kernel (see (1)). We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch [9] for surrogate training. To compare with other high-dimensional BO methods, we considered two training schemes. One performs MAP estimation with LBFGS. The maximum number of iterations was 15K and tolerance level 2E-9. We employed a diffuse prior over each length-scale, i.e., Uniform(0.001, 30). The noise variance was assigned a diffuse Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15). We denote this method as SBO-1. The second training scheme simply uses ADAM optimization [17] to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$. We denote this method as SBO-2. We also tested with Matérn-5/2 kernel with a single length-scale parameter, which we denoted as SBO-3. We used BOTorch [2] for acquisition function maximization, which uses LBFGS. Additional evaluations on alternative choices of the kernel, training scheme, and acquisition function are provided in Appendix Section B.
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [34], HESBO [25], ALEBO [18], SaasBO [6], and TURBO [7]. The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations. This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. We used the original implementation and default (recommended) settings for the competing methods. In addition, we compared with two BO methods based on Bayesian neural networks (BNN). One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference. We used a high-quality open-source implementation for each BNN based method. Table 3 summarizes the structural assumptions made by all the methods. The details of all the competing methods are given in Appendix Section A.2.
For each optimization task, we randomly queried the target function for 20 times to generate the initial data. We then ran each method for 400 steps for synthetic tasks and for 300 steps for real-world tasks. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For a reliable comparison, we ran every method for ten times.
4.2 Optimization Performance
Ultimate Outcomes. We first examined the ultimate optimization outcomes of each method. The results are reported in Fig. 1 for synthetic benchmarks and in Fig. 2 for real-world benchmarks. Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region). We can observe that on average, the standard BO, including both SBO-1 and SBO-2, outperforms the competing methods at every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NAS201, the performance of SBO is slight worse than (but still very close to) the best baseline(s).
It is noteworthy that the performance of the competing high-dimensional BO methods appears to be sensitive to the structure within the target function. For example, SaasBO imposes a strong inductive bias that only a small number of variables is related to the objective (via the usage of the horse-shoe prior). The target function in Hartman6(300, 6) and DNA aligns well with this assumption. As a result, SaasBO exhibits outstanding performance in both tasks. However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200) and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop. Another example is RDUCB and Tree-UCB, which assume that the target function can be decomposed as a summation of a group of functions, each over a small number of variables. RDUCB and Tree-UCB perform well in synthetic benchmarks that match this assumption, such as Rosenbrock(100, 100), Rosenbrock(300, 100), and StybTang(200, 200). However, their performance drops significantly in cases that do not align with their assumption, such as Ackley(150, 150) and Ackley(300, 150). Across all the benchmarks, the BNN based methods performed worse than most GP based methods. This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
In contrast, the standard BO exhibits much greater robustness to the structure of the target function. Using standard GP equipped with Matérn kernels, both SBO-1 and SBO-2 consistently rank among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc. Interestingly, as shown in Appendix Fig.7, even using a single length-scale parameter, the standard BO, i.e., SBO-3, still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-3 struggled in finding right locations to query, leading to poor performance.
Runtime Performance. Next, we examined the runtime performance. For each method, we report the average maximum function value obtained at each step in Fig. 5 and Fig. 6 of the Appendix. We supplement the runtime standard deviation and the result of SBO-3 in Fig. 8 of the Appendix. Note that since we have twelve baseline methods, reporting the standard deviations of all the methods makes the figure very cluttered. To ensure clarity, in Fig. 8, we show the standard deviation of the best, worst, and median performed baselines so far at each step.
One can see that after a few initial steps, SBO-1 and SBO-2 (and SBO-3 in most cases) consistently generated superior queries, finding larger function values. Their curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function. This trend is particularly evident in examples such as Rosenbrock(100, 100), StybTang(200, 200), Mopta08, Rover, SVM, and more.
The collective results demonstrate that the standard GP (with Matérn kernels) is effective for high-dimensional optimization. In comparison with more complex surrogate models using additional assumptions, the standard GP exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and enables superior optimization results.
Figure 1: Optimization performance in synthetic benchmarks. The results were summarized from 10 runs.
Figure 2: Optimization performance in real-world problems. The results were summarized from 10 runs.
4.3 Surrogate Learning
Next, we explicitly evaluated the performance of the standard GP in learning high-dimensional functions to confirm if it can serve as a reliable surrogate model for high-dimensional BO. To this end, from the optimization trajectories of all the methods, we randomly sampled 300 training examples from the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388). We sampled another 100 examples for testing. We compared with the surrogate models used in SaasBO, namely, placing a horse-shoe prior over the the length-scale parameters to prune the variables. We also compared with a random projection GP, denoted as ProjGP, to align with the baselines employing low-dimensional embeddings. ProjGP employs the kernel function in the form of $\vartheta({\bf x},{\bf x}^{\prime})=\kappa({\bf A}{\bf x},{\bf A}{\bf x}^{\prime})$, where ${\bf A}$ is a random projection matrix. We varied the rank of ${\bf A}$ from 10 and 20. Each element is sampled from the standard normal distribution. In addition, we compared with an additive GP, denoted as AddGP, where the kernel takes the form $\vartheta({\bf x},{\bf x}^{\prime})=\sum_{j}\kappa({\bf x}^{j},{\bf x}^{\prime j})$ where each ${\bf x}^{j}\subset{\bf x}$ and ${\bf x}^{\prime j}\subset{\bf x}^{\prime}$, ${\bf x}$ and ${\bf x}^{\prime}$ are the union of ${\bf x}^{j}$’s and ${\bf x}^{\prime j}$’s, respectively. AddGP imposes a decomposition structure in the functional space; see (5). We tested AddGP with randomly partitioning the input into 10 variables per group. We tested two versions: one using non-overlapping partitions, and the other overlapping partitions with one variable shared by successive groups. We trained the standard GP with ARD Matérn-5/2 kernel and with a single length-scale Matérn-5/2 kernel using MLE. SaasBO was trained by NUTS. For a fair comparison, ProjGP and AddGP were based on ARD Matérn-5/2 and trained via MLE as well.
(a) ARD Matérn
(b) Matérn
(c) SaasBO-NUTS
Figure 3: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row includes the results for Mopta08 (124) and the second row DNA (180).
We repeated the experiment for five times, and evaluated mean-square-error (MSE), the average test log-likelihood (LL), and the standard deviation. Table 4 and Appendix Table 5 report the results for all the GP models. The best and second best results are highlighted with a bold font. As we can see, in every dataset, standard GP with Mateŕ kernels achieves either the best or the second best result, demonstrating excellent and robust performance in learning high-dimensional functions.
In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180). As illustrated in Fig. 3, the prediction (and the predictive standard deviation) of ARD Matérn matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over Matérn (using a single length-scale parameter).
Together these results demonstrate that directly incorporating high-dimensional inputs into the kernel function does not necessarily compromise the learning performance. This implies that, even when the input space is high dimensional, the popular Matérn kernel remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling. On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to optimization performance. Our observation is consistent with the recent work in kernel operator learning [3], where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings. In this context, the input comprises of a discretized function, typically of very high-dimensionality (e.g., several thousands).
4.4 Computational Efficiency
So far, in our BO loop, the GP model is trained via MAP or MLE. Although the performance has been promising, it is worth investigating whether a full Bayesian approach would yield substantial improvements. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP with ARD Matérn. We adopted the same settings as used in SaasBO-NUTS [6]: The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 4 , and Appendix Fig. 9. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal. However, running NUTS is much more computationally costly compared to doing MAP and MLE. For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step. Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
By contrast, for more complex surrogate models, such as SaasBO, the MCMC based training might be more crucial for achieving good performance. As depicted in Fig. 2, SaasBO-NUTS demonstrates a substantial improvement over SaasBO-MAP across all real-world benchmarks, excluding NAS201. However, the computational cost of running SaasBO-NUTS is considerably higher, ranging from 5.4x (NAS201) to 33.9x (SVM, DNA) compared to the running time of SaasBO-MAP. The substantial increase in computational costs may pose a bottleneck for practical usage. Detailed per-optimization-step running time for these methods are provided in Appendix Table 6 and 7.
Dataset ARD Matérn Matérn SaasBO Proj-GP Add-GP
Mopta08(124) 0.034$\pm$0.006 0.145$\pm$0.057 0.038$\pm$0.008 0.461$\pm$0.157 0.107$\pm$0.033
Rover(60) 0.327$\pm$0.078 0.284$\pm$0.09 0.284$\pm$0.09 0.548$\pm$0.097 0.321$\pm$0.097
DNA(180) 0.149$\pm$0.038 0.332$\pm$0.053 0.152$\pm$0.049 0.747$\pm$0.121 0.312$\pm$0.061
SVM(388) 0.034$\pm$0.01 0.103$\pm$0.017 0.034$\pm$0.014 0.392$\pm$0.071 0.077$\pm$0.012
Ackley(150, 150) 0.017$\pm$0.008 0.007$\pm$0.001 0.007$\pm$0.001 0.034$\pm$0.007 0.007$\pm$0.001
Ackley150(300, 150) 0.023$\pm$0.006 0.012$\pm$0.001 0.011$\pm$0.001 0.080$\pm$0.024 0.021$\pm$0.003
Hartmann6 (300, 6) 0.017$\pm$0.013 0.419$\pm$0.048 0.024$\pm$0.023 0.583$\pm$0.094 0.416$\pm$0.084
Rosenbrock(100, 100) 0.084$\pm$0.013 0.118$\pm$0.023 0.124$\pm$0.025 0.378$\pm$0.069 0.097$\pm$0.015
Rosenbrock(300, 100) 0.090$\pm$0.064 0.160$\pm$0.09 0.131$\pm$0.075 0.533$\pm$0.214 0.138$\pm$0.075
Stybtang(200, 200) 0.043$\pm$0.018 0.036$\pm$0.014 0.039$\pm$0.012 0.253$\pm$0.040 0.038$\pm$0.009
Table 4: Mean Squared Error(MSE) of each surrogate model. The results were averaged from five runs.
Figure 4: Optimization performance when using ARD Matérn trained with MAP and MLE, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
4.5 Ablation Study and Failure Modes with Other Settings
Finally, we conducted a careful ablation study on the alternative settings to carry out standard BO. We found, however, those settings, such as using SE kernel and EI acquisition, often lead to inferior performance in high-dimensional optimization, which align with the a few failure cases reported in the literature [15, 20]. The detailed results and analysis are given in Section B of the Appendix. We summarize the favorable and unfavorable settings for high-dimensional optimization in Table 1.
5 Conclusion
We have presented a comprehensive investigation on standard BO (SBO) in high-dimensional optimization. We discovered that simply using Matérn kernel and UCB, SBO has already achieved the state-of-the-art performance. There seems no need for imposing additional structure assumptions for function learning and/or placing strong priors over the model parameters. In the future, we will continue applying SBO in more tasks, and seek for theoretical justifications and guarantees.
Appendix A More Experiment Details
A.1 Definition of Synthetic Functions
Stybtang. We used the following slightly-modified Stybtang function,
$f({\bf x})=\frac{1}{2}\sum_{i=1}^{D}\left((x_{i}-c_{i})^{4}-16(x_{i}-c_{i})^{2}+5(x_{i}-c_{i})\right),$
where ${\bf x}\in[-5,5]^{D}$, and we set $[c_{1},\dots,c_{D}]$ as an evenly spaced sequence in $[0,7.5]$ ($c_{1}=0$ and $c_{D}=7.5$). The optimum is at ${\bf x}^{\dagger}=[c_{1}-2.903534,\ldots,c_{D}-2.903534]$.
Rosenbrock. We used the following Rosenbrock function,
$f({\bf x})=\sum_{i=1}^{D-1}\left[100\left((x_{i+1}-c_{i+1})-(x_{i}-c_{i})^{2}\right)^{2}+\left(1-(x_{i}-c_{i})\right)^{2}\right],$
where ${\bf x}\in[-2.048,2.048]^{D}$. We set $[c_{1},...,c_{D}]$ as an evenly spaced sequence in $[-2,2]$, where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
Ackley. We used the Ackley function defined in [30],
$f({\bf x})=-20\text{ exp}\left(-0.2\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\text{exp}\left(\frac{1}{d}\sum_{i=1}^{d}\text{cos}(2{\pi}x_{i})\right)+20+\text{exp}(1),$
where each $x_{i}\in[-32.768,32.768]$ and the (global) optimum is at $x^{\dagger}=[0,\ldots,0]$.
Hartmann6. The function is given by
$f({\bf x})=-\sum_{i=1}^{4}{\alpha}_{i}{\exp}\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right),$
where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in [30]. The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
Figure 6: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
Figure 7: Ultimate optimization performance: Maximum Function Value obtained at the end of optimization loop, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel. The results were collect from 10 runs.
Figure 8: Runtime performance: Maximum Function Value obtained vs. Number of Steps in real-world benchmarks, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel. The results were averaged from 10 runs. The shaded region depicts the standard deviation.
A.2 Implementation
  • ALEBO. We used ALEBO implementation shared by the Adaptive Experimentation (AX) Platform (version 0.2.2). The source code is at https://github.com/facebook/Ax/blob/main/ax/models/torch/alebo.py.
  • HESBO. We used HESBO implementation of the original authors (https://github.com/aminnayebi/HesBO).
  • TURBO. We used the original TURBO implementation (https://github.com/uber-research/TuRBO)
  • SaasBO-NUTS. We used the original implementation of the authors (https://github.com/martinjankowiak/saasbo).
  • SaasBO-MAP. The SaasBO implementation available to the public does not include the version using MAP estimation. We therefore implemented this method based on the original paper [6]. All the hyperparameter settings follow exactly the same as the original paper.
  • RDUCB and Tree UCB. The implementation of both methods is publicly available at https://github.com/huawei-noah/HEBO/tree/master/RDUCB.
[…]
A.3 Computing Environment
We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips.
Task ARD Matérn Matérn SaasBO Proj-GP Add-GP
Mopta08(124) 0.570$\pm$0.08 -0.277$\pm$0.101 0.560$\pm$0.064 -0.962$\pm$0.183 -0.089$\pm$0.105
Rover(60) -0.975$\pm$0.216 -0.647$\pm$0.158 -0.638$\pm$0.14 -0.995$\pm$0.116 -0.703$\pm$0.125
DNA(180) -0.256$\pm$0.261 -0.666$\pm$0.102 -0.183$\pm$0.158 -1.137$\pm$0.143 -0.679$\pm$0.143
SVM(388) 0.581$\pm$0.119 -0.179$\pm$0.042 0.650$\pm$0.12 -0.837$\pm$0.095 0.062$\pm$0.129
Ackley(150, 150) 0.428$\pm$0.222 0.633$\pm$0.067 0.574$\pm$0.068 -0.006$\pm$0.268 0.569$\pm$0.066
Ackley(300, 150) 0.232$\pm$0.155 0.252$\pm$0.057 0.279$\pm$0.077 -0.372$\pm$0.185 0.220$\pm$0.074
Hartmann6(300, 6) 1.200$\pm$0.169 -0.625$\pm$0.079 1.528$\pm$0.564 -0.927$\pm$0.084 -0.617$\pm$0.118
Rosenbrock(100, 100) 0.640$\pm$0.081 0.445$\pm$0.062 0.526$\pm$0.08 -0.466$\pm$0.182 0.521$\pm$0.057
Rosenbrock(300, 100) 0.518$\pm$0.225 0.121$\pm$0.155 0.357$\pm$0.21 -0.706$\pm$0.209 0.189$\pm$0.2
Stybtang(200, 200) 0.574$\pm$0.164 0.563$\pm$0.073 0.543$\pm$0.067 -0.458$\pm$0.069 0.582$\pm$0.072
Table 5: Test log likelihood of each surrogate model. The results were averaged over five runs. Bolded are the best and second best results.
Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
ARD Matérn-MAP 7.39 9.99 7.81 9.77 13.55
ARD Matérn-MLE 11.23 9.84 5.76 14.34 32.16
ARD Matérn-NUTS 492.62 404.17 393.62 575.03 768.29
SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
Table 6: Average running time (in seconds) per optimization step in real-world benchmark tasks.
Rosenbrock Ackley Stybtang Rosenbrock Ackley Hartmann6
(100,100) (150, 150) (200, 200) (300, 100) (300, 150) (300, 6)
ARD Matérn-MAP 8.52 36.26 21.42 16.36 25.90 51.56
ARD Matérn-MLE 8.85 12.94 13.67 12.63 9.99 8.81
ARD Matérn-NUTS 857.21 850.23 1120.28 1276.42 1049.78 948.25
SaasBO-MAP 70.17 70.69 66.35 100.79 85.86 85.72
SaasBO-NUTS 1294.90 1621.60 1511.99 1938.53 925.80 2204.54
Table 7: Average running time (in seconds) per optimization step on synthetic benchmark tasks.
Figure 9: Optimization performance using MLE training, MAP training with a diffuse prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
Appendix B Ablation Study and Failure Modes With Other Settings
In this section, we conducted a careful ablation study to investigate alternative popular settings in executing standard BO. From these settings, we identified a few failure modes during high-dimensional optimization. We provided analysis and discussions accordingly.
B.1 GP Training
First, we examined MLE training with LBFGS. We found that, however, LBFGS often incurs numerical issues, and cannot reliably finish model training. This is consistent with the report in other works [12]. The issue might be because that without constraints (or regularization) over the parameters (such as the length-scales and amplitude), the exploration of LBFGS tends to be overly aggressive, which easily cause numerical problems. Hence, using stochastic optimization, like ADAM, might be a wise choice for MLE training.
Second, we examined MAP training with alternative prior choices over the length-scale parameters. In the main paper, we use a diffuse prior, Uniform(0.001, 30). In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular BO library BoTorch. To see how it is concentrated, in Fig. 10 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
We found that employing a concentrated prior over the length-scale parameters will severely deteriorate the high-dimensional optimization performance. As shown in Fig 12 and 13, after placing the prior Gamma(3,6) over each length-scale parameter, both the ultimate and runtime optimization performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training. To figure out the reason, we further investigated the surrogate learning performance. Following the same setting as in Fig. 3, we examined the prediction alignment after applying the prior Gamma(3,6) over the length-scale parameters. As depicted in Fig. 11 (third and fourth column), this concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function. This is consistent with the report in the SaasBO paper [6] (see Fig.1 of that paper). The quantitative results shown in Table 8 and 9 further confirm the substantially worse prediction accuracy as compared with GP training with a diffuse prior. In addition, we notice that the previous work[21] have also tested standard BO using Gamma(3,6) as the prior over the length-scale parameters. In their high-dimensional optimization experiments (see Fig. 5 of that paper), standard BO performs badly. We used our setting (SBO-1 and SBO-2) to re-run one experiment that maximizes a high-dimensional polynomial function. It turns out that both SBO-1 and SBO-2 can surpass the best method I-BNN reported by that paper.
Together these results have shown that a concentrated prior can severely limit GP’s capacity for high-dimensional function estimation, leading to inferior performance in high-dimensional optimization.
Figure 10: Probability density curves. Gamma(2, 0.15) and Gamma(1.1, 0.5) are the priors for the amplitude and noise variance, respectively.
Figure 11: Alignment between the prediction (y-axis) and ground-truth function values (x-axis).
Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
LBFGS ADAM
Mopta08(124) 0.034$\pm$0.006 0.035$\pm$0.005 0.837$\pm$0.112 0.856$\pm$0.114
Rover(60) 0.327$\pm$0.078 0.323$\pm$0.075 0.78$\pm$0.154 0.819$\pm$0.115
DNA(180) 0.149$\pm$0.038 0.166$\pm$0.049 1.014$\pm$0.097 1.026$\pm$0.101
SVM(388) 0.034$\pm$0.01 0.039$\pm$0.01 1.057$\pm$0.066 1.06$\pm$0.067
Ackley(150, 150) 0.017$\pm$0.008 0.026$\pm$0.007 0.019$\pm$0.004 0.019$\pm$0.005
Ackley(300, 150) 0.023$\pm$0.006 0.07$\pm$0.047 0.431$\pm$0.09 0.459$\pm$0.092
Hartmann6(300, 6) 0.017$\pm$0.013 0.178$\pm$0.325 0.881$\pm$0.045 0.906$\pm$0.034
Rosenbrock(100, 100) 0.084$\pm$0.013 0.081$\pm$0.012 0.9$\pm$0.156 0.925$\pm$0.172
Rosenbrock(300, 100) 0.09$\pm$0.064 0.116$\pm$0.078 0.971$\pm$0.222 1.001$\pm$0.217
Stybtang(200, 200) 0.043$\pm$0.018 0.054$\pm$0.024 0.698$\pm$0.05 0.732$\pm$0.034
Table 8: Mean Squared Error(MSE) averaged from five runs. The best results are bolded.
Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
LBFGS ADAM
Mopta08(124) 0.57$\pm$0.08 0.542$\pm$0.104 -1.337$\pm$0.058 -1.715$\pm$0.212
Rover(60) -0.975$\pm$0.216 -1.062$\pm$0.286 -$1$1.322$\pm$0.073
DNA(180) -0.256$\pm$0.261 -0.388$\pm$0.35 -1.426$\pm$0.054 -2.026$\pm$0.183
SVM(388) 0.581$\pm$0.119 0.504$\pm$0.128 -1.453$\pm$0.038 -2.095$\pm$0.129
Ackley(150, 150) 0.428$\pm$0.222 0.14$\pm$0.218 0.451$\pm$0.088 0.445$\pm$0.087
Ackley(300, 150) 0.232$\pm$0.155 -0.257$\pm$0.528 -1.006$\pm$0.091 -1.037$\pm$0.087
Hartmann6(300, 6) 1.2$\pm$0.169 0.67$\pm$0.867 -1.344$\pm$0.029 -1.618$\pm$0.128
Rosenbrock(100, 100) 0.64$\pm$0.081 0.606$\pm$0.1 -1.308$\pm$0.083 -1.598$\pm$0.42
Rosenbrock(300, 100) 0.518$\pm$0.225 0.347$\pm$0.325 -1.411$\pm$0.134 -1.967$\pm$0.413
Stybtang(200, 200) 0.574$\pm$0.164 0.421$\pm$0.228 -1.251$\pm$0.017 -1.41$\pm$0.067
Table 9: Test log likelihood averaged from five runs. The best results are bolded.
Figure 12: Ultimate optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters. SBO-1 and SBO-2 use Uniform(0.001, 30).
Figure 13: Runtime optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters. SBO-1 and SBO-2 use Uniform(0.001, 30).
B.2 Alternative Acquisition Functions
Next, we examined two other popular acquisition functions used in BO: Expected Improvement (EI) and Thompson sampling (TS). At each step, EI computes the expected improvement upon the best sample so far,
$\text{EI}({\bf x})=(\mu({\bf x})-f({\bf x}^{*}))\Psi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right)+\sqrt{v({\bf x})}\Phi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right),$ (6)
where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data. We used BOTorch to maximize EI through L-BFGS. To conduct TS at each step, we used Sobol Sequence [29] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior. We then selected the input with the largest function value as the next query. We denote this method by TS-3K. To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence. We denote this method as TS-10K. We used BOTorch implementation of both methods. We used ARD Matérn kernel and MLE training at each step.
The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig. 14 and 15, respectively. As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline. On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst.
The limitation of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1]. A remedy is to use log-EI as the acquisition function. The failure of the TS might be due to the high-dimension of the input space. A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function.
Figure 14: Ultimate performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
Figure 15: Runtime performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
B.3 Using Square Exponential (SE) kernels
Third, we considered another widely used kernel in BO, namely, the SE kernel. We tested with ARD SE and SE kernel using a single length-scale parameter. We used MLE training and UCB $(\lambda=1.5)$ to run BO. As depicted in Fig. 17 and 18, with smaller optimization dimensions, such as Mopta08 (124), NAS201 (3) and DNA (180), the performance of SE is among the best. However, with higher dimensions, e.g., above 200, the performance of SE drops substantially, such as in SVM (388) and Ackley(300, 150).
We argue that the inferior performance might be due to that the vanilla SE kernel is not powerful enough to enable GP to well capture high-dimensional functions. Consider a single length-scale parameter. we observe that the value of the SE kernel is determined by the squared distance between the inputs,
$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp\left(-\eta\|{\bf x}-{\bf x}^{\prime}\|^{2}\right).$ (7)
With the growth of the input dimension $d$, the kernel function value decays much faster than the increase of the distance: a squared decay coupling with the exponential decay. The learning can be easily driven to adapt to a sparse, diagonal-dominant covariance matrix, which roughly treats each data point as independent. This can make the learning fail to capture the correlation between function values, and result in inferior learning performance. By contrast, the value of the Matérn kernel (see (1)) is dominated by the distance rather than the squared distance. Therefore, with the growth of the distance, the kernel function value decays much slower than the SE kernel. It is easier for the optimization to adjust the length-scale parameter so as to avoid “diagonal-like” covariance matrices. Consequently, the correlation between function values can be better captured in the high dimensional input space.
To verify our argument, we examined the GP learning performance with ARD SE kernel and ARD Matérn kernel. We used Rosenbrock(d,d) and Hartmann6(d, 6) as test cases. The former is a “full-rank” function, namely, all the $d$ input variables are relevant to the function value. The latter has a “low-rank” structure in that only the first 6 variables out of the $d$ input variables determines the function value. For each test case, we increased the input dimension from 50 to 300. We randomly generated 500 training examples and 100 test examples. We then trained each method with MLE, and examined the mean square error (MSE) and the test log likelihood. We repeated the training and test procedure for five times, and report the average MSE, average test log likelihood, and the standard deviations in Fig. 16.
As we can see, in both test cases, when the input dimension $d$ does not exceed 150, the performance of using SE and Matérn-5/2 is close, in terms of both MSE and log likelihood. The two curves almost overlap. However, for $d>150$ in both test scenarios, the MSE significantly escalates when employing the SE kernel, accompanied by a substantial decrease in the test log likelihood. These observations suggest a failure of the SE kernel in effectively learning the target function. In comparison, employing the Matérn kernel maintains a consistently favorable prediction performance even when $d>150$. On the Rosenbrock dataset, the MSE gradually increases while the test log likelihood decreases, following a roughly linear trend. This outcome is expected as, with the rise in input dimension $d$, more training examples are necessary to sustain comparable performance. Furthermore, on the Hartmann6 dataset, little variation is observed in both MSE and test log likelihood with increasing the input dimension when using the Matérn kernel. This suggests that the GP has effectively discerned the pertinent input variables and mitigated the influence of irrelevant ones. In summary, these results collectively underscore the advantage of employing Matérn kernels for learning high-dimensional functions. It is worth noting that the most recent work [12] designs a log-normal prior over each length-scale parameter to overcome the learning issue with the ARD SE kernel.
Figure 16: Learning high-dimension functions with different kernels. We trained on 500 points, and tested on 100 points. Results shown with mean and standard deviation over 5 runs.
Figure 17: Ultimate optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
Figure 18: Runtime optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
B.4 Different Choice of $\lambda$ in UCB
Finally, we examined how the exploration parameter $\lambda$ in UCB (see (4)) influences the performance in high-dimensional optimization. To this end, we used ARD Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}. We conducted MLE training via ADAM optimization. We show the ultimate and runtime optimization performance in Fig. 19 and 20, respectively. It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close. However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$. This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
Figure 19: Ultimate optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.
Figure 20: Runtime optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.