### v4

# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

Abstract
A long-standing belief holds that Bayesian Optimization (BO) with standard Gaussian processes (GP) — referred to as standard BO — underperforms in high-dimensional optimization problems. While this belief seems plausible, it lacks both robust empirical evidence and theoretical justification. To address this gap, we present a systematic investigation. First, through a comprehensive evaluation across eleven widely used benchmarks, we found that while the popular Square Exponential (SE) kernel often leads to poor performance, using Matérn kernels enables standard BO to consistently achieve top-tier results, frequently surpassing methods specifically designed for high-dimensional optimization. Second, our theoretical analysis reveals that the SE kernel’s failure primarily stems from improper initialization of the length-scale parameters, which are commonly used in practice but can cause gradient vanishing in training. We provide a probabilistic bound to characterize this issue, showing that Matérn kernels are less susceptible and can robustly handle much higher dimensions. Third, we propose a simple robust initialization strategy that dramatically improves the performance of the SE kernel, bringing it close to state-of-the-art methods, without requiring any additional priors or regularization. We prove another probabilistic bound that demonstrates how the gradient vanishing issue can be effectively mitigated with our method. Our findings advocate for a re-evaluation of standard BO’s potential in high-dimensional settings.
1 Introduction
Many applications require optimizing complex functions, allowing queries only for the function values, possibly with noises, yet without any gradient information. Bayesian Optimization (BO) [27, 21] is a popular approach for addressing such challenges. BO typically employs Gaussian process (GP) [32] as a probabilistic surrogate. It iteratively approximates the target function, integrates the posterior information to maximize an acquisition function for generating new inputs at which to query, then updates the GP model with new examples, and concurrently approach the optimum.
Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems [8, 23, 6, 22, 18, 31, 19]. It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 23]. A continuous line of research is dedicated to developing novel high-dimensional BO methods. The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs. One class of methods assumes a decomposition structure within the functional space [16, 24, 10, 33], where the target function is expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables. Another family of methods assumes that the inputs are intrinsically low-rank. These methods either project the original inputs into a low-dimensional space [31, 23, 18] or use sparse-inducing priors to trim down massive input variables [6]. Subsequently, GP surrogates are built with the reduced input dimensions.
While the aforementioned concerns may seem plausible, there is a lack of both strong empirical evidence and theoretical justification to confirm and explain why standard BO would be ineffective in high-dimensional optimization. To bridge this gap, we systematically investigated standard BO in this paper. The major contributions of our work are summarized as follows:
  • Empirical Results. We investigated BO with standard GP across eleven widely-used benchmarks, including six synthetic and five real-world high-dimensional optimization tasks. The number of variables ranged from 30 to 388. We compared standard BO with nine state-of-the-art high-dimensional BO methods, and performed extensive evaluations and ablation studies. Surprisingly, while the popular SE kernel often led to poor performance, switching to the ARD Matérn kernel enabled standard BO to consistently achieve the best (or near-best) optimization performance across all the tasks. When successful, standard BO seems more flexible in accommodating various structures within the target functions.
  • Theory. Through analyzing the gradient structure of the GP likelihood, we identified the primary failure mode of standard BO, in particular when using the SE kernel. Specifically, the commonly used initializations for the length-scale parameters, such as setting them to one, is improper in high-dimensional settings and can easily cause gradient vanishing, preventing effective training. We proved a probabilistic tail bound to characterize this issue under mild conditions. Applying the same analytical framework to the Matérn kernel, we show through a comparison of the tail bounds that the Matérn kernel is less prone to gradient vanishing with the same length-scale initialization, making it more effective for handling high-dimensional problems.
  • Simple Robust Initialization. Based on our theoretical results, we proposed a simple yet robust length-scale initialization method, without requiring any additional priors or regularization. We proved a probabilistic bound showing that the probability of gradient vanishing decreases exponentially with the increase in dimensionality, implying that the gradient vanishing issue can be effectively mitigated in high-dimensional settings. Empirical evaluations demonstrate that our initialization method dramatically improves the performance of standard BO with SE kernels, enabling it to achieve, or come close to, the state-of-the-art performance in high-dimensional optimization.
2 Standard Bayesian Optimization
Consider maximizing a $d$-dimensional objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$. The form of $f(\cdot)$ is unknown. We can query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, but no gradient information is accessible. We aim to find ${\bf x}^{\dagger}=\mathrm{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$. To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to predict $f$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction. This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate. The utility typically encapsulates an exploration-exploitation trade-off. By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum. Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy. BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model. Then the iterative procedure repeats until convergence or a predefined stopping criterion is met.
The standard BO adopts GP regression [32] for surrogate modeling. A GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is usually chosen as a Mercer kernel function. One most popular kernel used in BO is the Square Exponential (SE) kernel with Automatic Relevance Determination (ARD),
$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp(-\rho^{2}),$ (1)
where $a>0$ is the amplitude, $\rho=\sqrt{({\bf x}-{\bf x}^{\prime})^{\top}{\rm diag}(\frac{1}{\bm{\ell}^{2}})({\bf x}-{\bf x}^{\prime})}$, and $\bm{\ell}=(\ell_{1},\ldots,\ell_{d})^{\top}>{\bf 0}$ are the length-scale parameters. We refer to this kernel as ARD because each input dimension has a distinct length-scale parameter. An alternative choice is the ARD Matérn-5/2 kernel,
[…]
Given training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$, let us denote the function values at ${\bf X}$ as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, which according to the GP prior, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Then one can employ a Gaussian likelihood for the observations ${\bf y}$, and obtain
$p({\bf y}|{\bf X})=\mathcal{N}({\bf y}|{\bf m},{\bf K}+\sigma^{2}{\bf I}),$ (3)
where $\sigma^{2}$ is the noise variance. To estimate the GP parameters, e.g., the length-scales $\bm{\ell}$, one can maximize the marginal likelihood or use MCMC sampling. The predictive distribution (given the GP parameters) at any new input ${\bf x}^{*}$, is a closed-form, conditional Gaussian distribution, $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right)$, where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
In each iteration, given the current GP surrogate model, an acquisition function is maximized to identify the next input location at which to query. One commonly used acquisition function is the upper confidence bound (UCB), defined as $\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X}$, where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) [15], Thompson sampling (TS) [25], the recently proposed log-EI [1], among others.
3 High Dimensional Bayesian Optimization
An enduring and widespread belief is that when the input dimension $d$ is high, e.g., a few hundreds, the standard BO is prone to failure. This belief might partly arise from an intuition that commonly used kernels, such as (1), could encounter challenges in dealing with high-dimensional inputs, making GP struggle in capturing the target function. A dedicated line of research is developing novel high-dimensional BO methods. The key idea of these methods is to impose some structural assumptions in surrogate modeling to sidestep directly handling the high-dimensional inputs in kernels.
Structural Assumption in Functional Space. The first class of methods assumes a decomposition structure within the functional space. Specifically, the target function $f$ is modeled as
$f({\bf x})=\sum\nolimits_{j=1}^{M}f_{j}({\bf x}^{j}),\;\;f_{j}\sim\mathcal{GP}\left(m_{j}({\bf x}^{j}),\kappa_{j}(\cdot,\cdot)\right),$ (4)
where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$. There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$. In [16], the input ${\bf x}$ is partitioned into non-overlapped groups. After every a few BO iterations, a new partition is selected from a set of random partitions. The selection is based on the model evidence. In [24], the variable groups can overlap, and are represented as the maximum cliques on a dependency graph. A Gibbs sampling method was developed to learn the structure of the dependency graph. Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning. However, the more recent work [33] points out that, learning the group structure through relatively small data can be misleading. A wrongly learned structure can cause queries to be stuck at local optima. The work proposes RDUCB that randomly decomposes ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
Structural Assumption in Input Space. Another class of methods assumes a low-rank structure in the input space. Many of them introduce Low-dimensional Embeddings. Wang et al., [31] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al., [23] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially imposes further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  [18]. ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue. Another direction is to impose Partial Variable Dependency by triming down the input variables. The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [12]. While one can also use MAP estimation, the performance is often much worse.
4 Theoretical Analysis
To understand why standard BO can fail in high-dimensional cases, in particular with the SE kernel (see Section 6), we look into the gradient structure during the GP training. Specifically, we consider at the beginning, all the length-scale parameters are set to the same initial value, $\ell_{1}=\ldots=\ell_{d}=\ell_{0}$. From the marginal likelihood (3), we can derive the gradient w.r.t each length-scale,
$\frac{\partial\log p({\bf y}|{\bf X})}{\partial\ell_{k}}=\frac{1}{2}{\rm tr}\left({\bf A}\cdot\frac{\partial{\bf K}}{\partial\ell_{k}}\right),$ (5)
where ${\bf A}=\bm{\alpha}\bm{\alpha}^{\top}-({\bf K}+\sigma^{2}{\bf I})^{-1}$, and $\bm{\alpha}=({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$. When using the SE kernel as specified in (1), each entry of $\frac{\partial{\bf K}}{\partial\ell_{k}}$ is as follows: $\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ii}=0$ and for $i\neq j$,
$\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ij}=\frac{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})}{\partial\ell_{k}}=\frac{2a}{\ell_{k}}\exp(-\frac{\|{\bf x}_{i}-{\bf x}_{j}\|^{2}}{\ell_{0}^{2}})\frac{(x_{ik}-x_{jk})^{2}}{\ell_{k}^{2}}\leq\frac{2a}{\ell_{k}}\cdot\frac{\rho^{2}}{e^{\rho^{2}}},$ (6)
where $x_{ik}$ and $x_{jk}$ are the $k$-th element in ${\bf x}_{i}$ and ${\bf x}_{j}$, respectively, and $\rho=\frac{\|{\bf x}_{i}-{\bf x}_{j}\|}{\ell_{0}}$.
Intuitively, as the input dimension $d$ increases, the squared distance between input vectors grows rapidly. Consequently, the factor $\frac{\rho^{2}}{e^{\rho^{2}}}$ in (6) can quickly fall below the machine epsilon $\xi$, causing the gradient vanishing issue [3, 11]. That is, the length-scale cannot be effectively updated according to $\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}$, due to the limitations of the numerical precision.
Proposition 4.1.
Given any $\xi>0$, $\frac{\rho^{2}}{e^{\rho^{2}}}<\xi$ when $\rho>\tau_{\text{SE}}=\frac{1}{2}+\sqrt{\frac{1}{4}-\log\xi}$.
When using the double-precision floating-point format (float64), the rounding machine epsilon is $\xi=2^{-53}$, and the above threshold is $\tau_{\text{SE}}=6.58$. To analyze the likelihood of gradient vanishing, we provide a probabilistic bound under uniform data distribution.
Lemma 4.2.
Suppose the input domain is $[0,1]^{d}$ and the input vectors are sampled independently from the uniform distribution, then for any constant threshold $\tau>0$, when $d>6\ell_{0}^{2}\tau^{2}$,
$p(\rho\geq\tau)>1-2\exp\left(-\frac{(d-6\ell_{0}^{2}\tau^{2})^{2}}{18d}\right).$ (7)
The lower bound grows exponentially with the input dimension $d$ and converges to one. This implies that given any fixed choice of $\ell_{0}$ relevant to $d$ (e.g., $\ell_{0}=1$), as $d$ increases, the probability $p(\rho\geq\tau_{\text{SE}})$ will rapidly approach one, leading to vanishing gradient for every $\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})$. As a consequence, each length scale $\ell_{k}$ in (5) cannot be effectively updated, preventing meaningful training [^1] A union bound can be directly derived from (7) to lower bound the joint probability that all gradients $\{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}\}_{i,j}$ across $N$ training instances are below machine epsilon. However, this bound is too loose to reflect our practical observation (see Fig. 1 and  2 and Table 2). Therefore, we focus our discussion on a single pair of inputs. . The inferior prediction accuracy of the GP model further contributes to poor performance of the BO procedure. Since $\tau_{\text{SE}}$ is small, gradient vanishing can occur even when $d$ is not large. For example, for $\ell_{0}=0.5$ and $\ell_{0}=1$, we have $p(\rho>\tau_{\text{SE}})>0.99$ when $d\geq 205$ and $d\geq 473$, respectively.
Why Matérn kernels perform much better? In our practical evaluations (see Section 6), standard BO with Matérn kernels (2) often perform much better than with the SE kernel in high-dimensional problems, frequently achieving state-of-the-art results. To understand the reasons behind this improvement, we apply the same framework to analyze the behavior of GP training with the Matérn kernel. First, we obtain the gradient,
[…]
As the distance increases, the factor $\rho^{2}/e^{\sqrt{5}\rho}$ will also converge to zero, implying that with the growth of $d$, the gradient will eventually fall below machine epsilon. However, since this factor $\rho^{2}/e^{\sqrt{5}\rho}$ decreases much slower than the corresponding factor to the SE kernel, namely $\rho^{2}/e^{\rho^{2}}$ in (6), the Matérn kernel is able to robustly handle much higher dimensions.
Proposition 4.3.
[…]
Lower bound 0.95 0.99 0.995 0.999 0.9995 0.9999
SE ($d$) 172 205 219 250 264 294
Matérn ($d$) 980 1040 1064 1116 1137 1185
Table 1: The input dimension $d$ vs. lower bound probability of gradient vanishing under $\ell_{0}=0.5$.
Numerical Verification. To verify whether gradient vanishing occurs during practical GP training and to access whether our analysis aligns with the empirical observations, we conducted numerical experiments using two synthetic functions, Hartmann6($d$, 6) and Rosenbrock(d,d) (see Section 6.1 for definitions). We varied the input dimension $d$ across {50, 100, 200, 300, 400, 500, 600}, with the domain set to $[0,1]^{d}$. For each value of $d$, we uniformly sampled $500$ training inputs and $100$ test inputs. We trained GP with both SE and Matérn kernels, using length-scale initialization $\ell_{0}$ from $\{0.1,0.5,0.693,1.0,\sqrt{d}\}$. Note $\ell_{0}=0.693=\text{SoftPlus}(0)$ is a popular initialization choice, used as the default choice in influential GP/BO libraries GPyTorch [9] and BOTorch [2]. Training and testing were repeated 20 times for each $d$ and $\ell_{0}$. We then evaluated (1) the average relative $L_{2}$ difference between the length-scale vector before and after training, namely, $\|\bm{\ell}_{\text{trained}}-\bm{\ell}_{\text{init}}\|/\|\bm{\ell}_{\text{init}}\|$, (2) the average gradient norm at the first training step, and (3) the average test Mean-Squared-Error (MSE).
As shown in Fig. 1, gradient vanishing indeed occurred, leading to training failure. When using $\ell_{0}=0.1$ for GP training with the SE and Matérn kernels, both the length-scale relative $L_{2}$ difference and the gradient norm remained close to zero across all choices of $d$ (those values are not displayed because they are too small). Consequently, the MSE hovered around 1.0, indicating consistent training failure. When we increased $\ell_{0}$ to 0.5 and 1.0 for training with the SE kernel, both the length-scale difference and gradient norm became significantly larger and non-trivial for $d=50$ and $d=100$, resulting in MSE below 0.2. This demonstrates that gradient vanishing did not occur and training was successful. However, as $d$ increased, gradient vanishing reappeared at $d=200$ for $\ell_{0}=0.5$ and at $d=400$ for $\ell_{0}=1.0$, for both Hartmann6 and Rosenbrock. In these cases, the length-scale difference and gradient norm dropped near zero, while the MSE returned to around 1.0. A similar trend was observed with the Matérn kernel. However, GP training with the Matérn kernel can succeed with higher dimensions under the same initialization conditions. For example, with $\ell_{0}=0.5$, for both Hartmann6 and Rosenbrock, training with the Matérn kernel started to failed until $d=400$, compared to $d=200$ with the SE kernel. When using $\ell_{0}=1.0$, the Matérn kernel succeeded across all $d$ values, while the SE kernel started to fail at $d=400$. Table 2 summarizes the input dimensions at which gradient vanishing and training failure began to happen.
Confirmation in BO Running. We next investigated whether gradient vanishing occurs during the BO process, potentially leading to poor optimization performance. To this end, we examined two benchmark BO tasks: SVM (D=388) and Rosenbrock (300,100), as detailed in Section 6. As shown in Fig. 3 and 4, standard BO with the Matérn kernel (SBO-Matérn) achieved top-tier results, while BO with the SE kernel (SBO-SE) performed poorly. Both methods used the initialization $\ell_{0}=0.693$. We then extracted the BO trajectories for both methods and analyzed the relative $L_{2}$ difference in the length-scale vectors before and after training, and the gradient norm at the first training step. As illustrated in Fig. 2, during the BO iterations, GP-SE ($\ell_{0}=0.693$) consistently experienced gradient vanishing and so training failure, whereas GP-Matérn ($\ell_{0}=0.693$) maintained non-trivial gradients. This difference directly correlates with their respective optimization performances — poor for SBO-SE and excellent for SBO-Matérn.
Figure 1: Mean Square Error (MSE), Relative $L_{2}$ difference between the length-scale vectors before and after training, and $L_{2}$ norm of the length-scale gradient at the first training step, across different input dimensions ($x$-axis) and initializations (legend). The results were averaged from 20 runs. For $\ell_{0}=0.1$, the length-scale relative difference and gradient norm are not displayed, as they fall much below the minimum values on the $y$-axis.
Figure 2: The relative $L_{2}$ difference of the length-scale vectors before and after training and the gradient norm at the first training iteration across every step of BO. The results for GP-SE ($\ell_{0}=0.693$) are not displayed, as they fall much below the minimum values on the y-axis.
Hartmann6 $\ell_{0}=0.1$ 0.5 0.6931 1.0 $\sqrt{d}$
SE 50 200 200 400
Matérn 50 400 600
Rosenbrock $\ell_{0}=0.1$ 0.5 0.6931 1.0 $\sqrt{d}$
SE 50 200 200 400
Matérn 50 400 600
Table 2: The input dimension from which gradient vanishing started to occur and training started to fail; ✓means no gradient vanishing occurred.
5 Robust Initialization
From the analysis in Section 4, one straightforward approach to improving learning with the SE (and also Matérn) kernel is to initialize each length-scale with a larger $\ell_{0}$, which can loosen the lower bound in (7) and allow the GP model to better adapt to higher dimensions. However, this approach is not robust enough, as $d-6\ell_{0}^{2}\tau^{2}$ will continue to increase with $d$, and the bound will still rise toward one rapidly, causing the gradient to vanish. To address this issue, we propose setting $\ell_{0}$ to growing with $d$ rather than keeping it constant,
$\ell_{0}=c\sqrt{d},\;\;\;c>0.$ (8)
Lemma 5.1.
Suppose the input domain is $[0,1]^{d}$ and each input vector is independently sampled from the uniform distribution. Given any constant threshold $\tau>0$, we set $\ell_{0}=c\sqrt{d}$ such that $c>\frac{1}{\sqrt{6}\tau}$, then
$p(\rho\geq\tau)\leq 2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right)\propto\exp(-\mathcal{O}(d)).$ (9)
With this initialization, the increase in $d$ exponentially reduces the upper bound on the probability of $\rho$ exceeding the given threshold $\tau$. As a result, the gradient vanishing issue can be fundamentally mitigated in the high-dimensional settings.
In our experiments, we tested standard BO using SE kernel with our proposed length-scale initialization as specified in (8) where we set $c=1$. As shown in Fig. 3 and Fig. 4, the performance across all the benchmarks is dramatically improved, as compared to using popular initialization $\ell_{0}=0.693$. The performance matches that of using Matérn kernels, achieving state-of-the-art (or near-best) results. In addition, as shown in Fig. 1 and Table 2, our numerical experiments confirm that gradient vanishing never occurred with our robust initialization method. Fig. 2 further demonstrates that our method eliminates gradient vanishing during practical BO runs, enabling effective GP training and much better optimization results. Together this shows that our method can effectively adapt SE kernels to high-dimensional settings, and mitigate the failure mode of standard BO.
Alternative method. Recent work [13] proposes an intuitive approach to address the learning challenge of the SE kernel in the high-dimensional spaces by constructing a log-normal prior over each length scale, $p(\ell_{k})=\text{LogNormal}(\ell_{k}|\mu_{0}+\frac{\log(d)}{2},\sigma_{0})$. This prior regularizes the length-scale to be on the order of $\sqrt{d}$ during training. Using our theoretical framework, this method can be justified as an alternative strategy to alleviate the gradient vanishing issue. However, our method, directly motivated by our analysis, is even simpler — requiring no additional prior construction or associated parameters, and has also shown effective in addressing the gradient vanishing issue. The comparison with  [13] is given in Section 6.
6 Comprehensive Evaluation
6.1 Experimental Settings
For a comprehensive evaluatio of the standard BO, we employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity, all the tasks aim for function maximization.
$\text{Fun}(d,d^{\prime})$ Structure
Ackley(300, 150) Partial variable dependency
Rosenbrok(300, 100) Partial variable dependency
Nonoverlap additive decomposition
Hartmann6(300,6) Partial variable dependency
Rosenbrock(100, 100) Nonoverlap additive decomposition
Stybtang(200, 200)
Ackley(150, 150) None
Table 3: Synthetic benchmarks and the structures within the target function: $d$ is the input dimension, and $d^{\prime}$ is the number of effective variables to compute the target function.
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section B.1 of the Appendix. Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The ground-truth of the target function values is computed with the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 3.
Real-World Benchmarks. We employed the following real-world benchmark problems: Mopta08 (124)  [14], a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints.
We followed [6] to encode the constraints as a soft penalty, which is added into the objective. SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model. Rover (60), a rover trajectory planning problem from [30]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment. DNA (180) [26], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [20]. Prior analysis [26] shows that only around 43 variables are relevant to the target function. NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
Methods. We tested with three versions of standard BO:
  • SBO-Matérn: GP regression with ARD Matérn-5/2 kernel as specified in (2), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$, as the default choice in the influential GP/BO libraries GPyTorch and BOTorch.
  • SBO-SE: GP regression with ARD SE kernel as specified in (1), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$.
  • SBO-SE(RI): GP regression with ARD SE kernel, using our proposed robust initialization $\ell_{0}=c\sqrt{d}$ where $d$ is the input dimension and $c=1$
We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch for GP training. To expose the possible gradient vanishing issue , we used Uniform(0.001, 30) over each length-scale. The noise variance was assigned a diffused Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15). We ran L-BFGS to obtain MAP estimation. The maximum number of iterations was 15K and tolerance level 2E-9.
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [33], HESBO [23], ALEBO [18], SaasBO [6], and TURBO [7]. The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations. This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. In addition, we tested the most recent method VBO (Vanilla BO[13] that levels up the performance of the standard BO by constructing a log-normal prior over the length-scale parameters (see Section 5). We used the original implementation and default (recommended) settings for the competing methods. Besides, we compared with two BO methods based on Bayesian neural networks (BNN). One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference. We used a high-quality open-source implementation for each BNN based method. Appendix Table 6 summarizes the structural assumptions made by all the methods. The details of all the competing methods are given in Appendix Section B.2.
For each optimization task, we randomly queried the target function for 20 times to collect the initial data. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For a reliable comparison, we ran every method for ten times.
6.2 Optimization Performance
Ultimate Outcomes. We first examined the ultimate optimization outcomes of each method. The results are reported in Fig. 3 for synthetic benchmarks and in Fig. 4 for real-world benchmarks. Due to space limit, we supplement the BNN-based BO results in Appendix Section C.1. Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
On average, SBO-Matérn, SBO-SE(RI) and VBO consistently deliver top-tier results, often outperforming the other methods by a large margin, e.g., on benchmarks Ackley (300, 150), DNA, and SVM. It is noteworthy that SBO-SE also performed exceptionally well when the input dimension $d\leq 180$, as seen in benchmarks Rosenbrock (100, 100), Mopta08, NAS, and DNA. However, as $d$ increases, e.g., Ackley (300, 150), Hartman6 (300, 6), Rosenbrock (300, 100), StybTang(200, 200) and SVM (388), the performance of SBO-SE dropped dramatically due to the gradient vanishing issue (see our theoretical analysis and numerical experiments in Section 4). After applying our robust initialization, the performance of SE kernels, namely, SBO-SE (RI) immediately recovered to top-tier levels, closely matching both SBO-Matérn and VBO. On the other hand, though SBO-Matérn used a small length-scale initialization ($\ell_{0}=0.693$), the Matérn kernel’s structure has made it far less susceptible to the gradient vanishing issue, enabling it to handle much higher dimensions than the SE kernel. Free from this issue, SBO-Matérn consistently achieved top-tier results.
Runtime Performance. Next, we examined the runtime performance. Fig. 5 reports the average maximum function value obtained at each step and the standard deviation across all runs. Since we have 16 methods for comparison, showing the standard deviation for every method makes the figure cluttered. To ensure clarify, we show the curves and standard deviation of SBO, VBO, and the best, worst, and median performed remaining baselines so far at each step.
After an initial stage, SBO-Matérn, SBO-SE (RI), and VBO consistently produced superior queries, achieving larger function values. Their performance curves are generally above those of competing methods, reflecting a more efficient optimization of the target function. This trend is particularly noticeable in benchmarks such as Ackley (300, 150), Rosenbrock (300, 100), and SVM. When $d>180$, however, the performance of SBO-SE declines, with its curve positioned toward the bottom due to the gradient vanishing issue. In benchmarks such as Rosenbrock (100, 100) and DNA (180), the curve for VBO is predominantly above those of SBO-Matérn and SBO-SE (RI), indicating even more efficient optimization. However, on Ackley (150, 150) and Ackley (300, 150), VBO’s curve is mostly below that of SBO-Matérn and SBO-SE (RI). For the remaining benchmarks, their curves are close and often overlap. Overall, in all cases, they tend to converge to similar function values.
These results collectively demonstrate that standard BO can excel in high-dimensional optimization problems. The lack of additional structural assumptions in GP modeling may make it more flexible in capturing various correlations within high-dimensional spaces. The primary challenge for standard BO arises from the gradient vanishing issue, which is often due to improper initialization of the length-scales. By using our robust initialization method, this risk can be significantly mitigated.
Additional Results. We conducted extensive additional evaluations and ablation studies to examine the impact of alternative GP training methods, exploration parameters in UCB, and other acquisition functions on SBO performance in high-dimensional settings. We found that MLE achieves optimization results comparable to the MAP training employed in our evaluation. However, using concentrated priors on the length-scale parameters (concentrated on small mean values) — such as the popular Gamma(3,6) prior that is the default choice in BOTorch — often leads to training failure and poor performance, aligning with our theoretical analysis. Both MLE and MAP perform similarly to Markov Chain Monte Carlo (MCMC) training in terms of optimization outcomes, but MCMC is significantly slower. UCB is notably robust to variations in the exploration parameter. Additionally, log-EI facilitates strong performance in high-dimensional optimization, whereas EI and TS often lead to subpar results. Detailed results and discussions are provided in Appendix Section C.
Figure 3: Optimization performance in synthetic benchmarks. The results were summarized from 10 runs.
Figure 4: Optimization performance in real-world problems. The results were summarized from 10 runs.
7 Conclusion
We conducted a thorough investigation of standard BO in high-dimensional optimization, both empirically and theoretically. Our analysis identified gradient vanishing as a major failure mode of standard BO, and we proposed a simple yet robust initialization method to address this issue. Our empirical evaluation demonstrates that, once this problem is mitigated, standard BO can achieve state-of-the-art performance in high-dimensional optimization.
Appendix A Proofs
A.1 Proof of Proposition 4.1
Proof.
We intend to solve $\rho^{2}/e^{\rho^{2}}<\xi$, which is equivalent to solving
$\log r-r<\log\xi,$
where $r=\rho^{2}$. We then leverage the fact that $\log r<\sqrt{r}$ when $r>1$. Since $1/e>\xi$, we know the solution of $\rho$ must be larger than 1, implying that the solution for $r$ is larger than $1$. To meet the inequality, we can solve
$\log r-r<\sqrt{r}-r<\log\xi.$
It then converts to a quadratic form,
$\sqrt{r}-\sqrt{r}^{2}<\log\xi.$ (10)
Note that $\rho=\sqrt{r}$. It is straightforward to solve this inequality and we can obtain that the inequality holds when
$\rho>\tau_{\text{SE}}=\frac{1}{2}+\sqrt{\frac{1}{4}-\log\xi}.$
A.2 Proof of Lemma 4.2
Proof.
First, since each $x_{ik}$ and $x_{jk}$ independently follow $\text{Uniform}(0,1)$, it is straightforward to obtain $\mathbb{E}[(x_{ik}-x_{jk})^{2}]=\frac{1}{6}$. Let us define $\gamma=\|{\bf x}_{i}-{\bf x}_{j}\|^{2}$. Then we have $\mathbb{E}[\gamma]=\frac{1}{6}d$. According to Hoeffding’s inequality, for all $t>0$,
$p(|\gamma-\frac{1}{6}d|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{d}\right).$ (11)
Since $\rho=\frac{\sqrt{\gamma}}{\ell_{0}}$, we have $p(\rho>\tau)=p(\gamma>\tau^{2}\ell_{0}^{2})$. Since $d>6\ell_{0}^{2}\tau^{2}$, we set $t=\frac{1}{6}d-\ell_{0}^{2}\tau^{2}>0$, and apply (11) to obtain
$p(|\gamma-\frac{1}{6}d|\geq\frac{1}{6}d-\ell_{0}^{2}\tau^{2}|)\leq 2\exp\left(-\frac{2\left(\frac{1}{6}d-\ell_{0}^{2}\tau^{2}\right)^{2}}{d}\right)=2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right),$
which is equivalent to
$p(|\gamma-\frac{1}{6}d|<\frac{1}{6}d-\ell_{0}^{2}\tau^{2})>1-2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right).$ (12)
Since $p(\gamma>\ell_{0}^{2}\tau^{2})\geq p(\ell_{0}^{2}\tau^{2}<\gamma<\frac{1}{3}d-\ell_{0}^{2}\tau^{2})=p(|\gamma-\frac{1}{6}d|<\frac{1}{6}d-\ell_{0}^{2}\tau^{2})$, combining with (12) we obtain the bound
$p(\rho>\tau)=p(\gamma>\ell_{0}^{2}\tau^{2})>1-2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right).$
A.3 Proof of Proposition 4.3
Proof.
To solve $\rho^{2}/e^{\sqrt{5}\rho}<\xi$, it is equivalent to solving
$\frac{1}{5}\frac{r^{2}}{e^{r}}<\xi,\;\;\;r=\sqrt{5}\rho,$
which is further equivalent to
$\log\frac{1}{5}+2\log r-r<\log\xi.$
We leverage $\log r<\sqrt{r}$ for $r>1$, and solve the upper bound,
$\log\frac{1}{5}+2\sqrt{r}-r<\log\xi,$
which gives
[…]
A.4 Proof of Lemma 5.1
Proof.
First, similar to the proof in Section A.2, under the uniform distribution on $[0,1]^{d}$, from Hoeffding’s inequality, we have for all $t>0$,
$p(|\gamma-\frac{1}{6}d|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{d}\right).$ (14)
When we set $\ell_{0}=c\sqrt{d}$ such that $c>1/\sqrt{6}\tau$, we have $\ell_{0}^{2}\tau^{2}>d/6$. Let us choose $t=\ell_{0}^{2}\tau^{2}-d/6=c^{2}\tau^{2}d-d/6>0$, and apply the Hoeffding’s inequality (14),
$p\left(|\gamma-d/6|\geq\ell_{0}^{2}\tau^{2}-d/6\right)\leq 2\exp\left(-\frac{2(\ell_{0}^{2}\tau^{2}-d/6)^{2}}{d}\right)=2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right).$ (15)
Since $p(\rho\geq\tau)=p(\gamma\geq\ell_{0}^{2}\tau^{2})\leq p(|\gamma-d/6|\geq\ell_{0}^{2}\tau^{2}-d/6)$, we obtain
$p(\rho\geq\tau)\leq 2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right).$
Appendix B More Experiment Details
B.1 Definition of Synthetic Functions
Stybtang. We used the following slightly-modified Stybtang function,
$f({\bf x})=\frac{1}{2}\sum_{i=1}^{D}\left((x_{i}-c_{i})^{4}-16(x_{i}-c_{i})^{2}+5(x_{i}-c_{i})\right),$
where ${\bf x}\in[-5,5]^{D}$, and we set $[c_{1},\dots,c_{D}]$ as an evenly spaced sequence in $[0,7.5]$ ($c_{1}=0$ and $c_{D}=7.5$). The optimum is at ${\bf x}^{\dagger}=[c_{1}-2.903534,\ldots,c_{D}-2.903534]$.
Rosenbrock. We used the following Rosenbrock function,
$f({\bf x})=\sum_{i=1}^{D-1}\left[100\left((x_{i+1}-c_{i+1})-(x_{i}-c_{i})^{2}\right)^{2}+\left(1-(x_{i}-c_{i})\right)^{2}\right],$
where ${\bf x}\in[-2.048,2.048]^{D}$. We set $[c_{1},...,c_{D}]$ as an evenly spaced sequence in $[-2,2]$, where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
Ackley. We used the Ackley function defined in [29],
$f({\bf x})=-20\text{ exp}\left(-0.2\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\text{exp}\left(\frac{1}{d}\sum_{i=1}^{d}\text{cos}(2{\pi}x_{i})\right)+20+\text{exp}(1),$
where each $x_{i}\in[-32.768,32.768]$ and the (global) optimum is at $x^{\dagger}=[0,\ldots,0]$.
Hartmann6. The function is given by
$f({\bf x})=-\sum_{i=1}^{4}{\alpha}_{i}{\exp}\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right),$
where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in [29]. The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in all real-world and synthetic benchmarks. The results were averaged from 10 runs. For cleaner view, we plotted the full results only for SBO and VBO. For the remaining methods, we show best, median, and worst result at each step.
Figure 6: Runtime performance comparison with BNN-based BO methods and SBO with a single length-scale parameter. Matérn kernel was used to conduct SBO. For clarity, VBO was excluded for comparison. The results were averaged from 10 runs. The shaded region depicts the standard deviation.
B.2 Implementation
  • ALEBO. We used ALEBO implementation shared by the Adaptive Experimentation (AX) Platform (version 0.2.2). The source code is at https://github.com/facebook/Ax/blob/main/ax/models/torch/alebo.py.
  • HESBO. We used HESBO implementation of the original authors (https://github.com/aminnayebi/HesBO).
  • TURBO. We used the original TURBO implementation (https://github.com/uber-research/TuRBO)
  • SaasBO-NUTS. We used the original implementation of the authors (https://github.com/martinjankowiak/saasbo).
  • SaasBO-MAP. The SaasBO implementation available to the public does not include the version using MAP estimation. We therefore implemented this method based on the original paper [6]. All the hyperparameter settings follow exactly the same as the original paper.
  • RDUCB and Tree UCB. The implementation of both methods is publicly available at https://github.com/huawei-noah/HEBO/tree/master/RDUCB.
[…]
B.3 Computing Environment
We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips.
Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
ARD Matérn-MAP 7.39 9.99 7.81 9.77 13.55
ARD Matérn-MLE 11.23 9.84 5.76 14.34 32.16
ARD Matérn-NUTS 492.62 404.17 393.62 575.03 768.29
SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
Table 4: Average running time (in seconds) per optimization step in real-world benchmark tasks.
Rosenbrock Ackley Stybtang Rosenbrock Ackley Hartmann6
(100,100) (150, 150) (200, 200) (300, 100) (300, 150) (300, 6)
ARD Matérn-MAP 8.52 36.26 21.42 16.36 25.90 51.56
ARD Matérn-MLE 8.85 12.94 13.67 12.63 9.99 8.81
ARD Matérn-NUTS 857.21 850.23 1120.28 1276.42 1049.78 948.25
SaasBO-MAP 70.17 70.69 66.35 100.79 85.86 85.72
SaasBO-NUTS 1294.90 1621.60 1511.99 1938.53 925.80 2204.54
Table 5: Average running time (in seconds) per optimization step on synthetic benchmark tasks.
Method Structural Assumption
ALEBO, HESBO Low-dim embedding
SaasBO Partial variable dependency
RDUCB, Tree UCB Additive function decomposition
SBO, TURBO, BNN None
Table 6: BO methods and their structural assumption about the target function.
Figure 7: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
Figure 8: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
Appendix C Additional Results
C.1 BNN-based BO and Standard BO with a Single Length-Scale
Due to the space limit, we did not report the results of the BNN-based BO in the main paper. Here, we supplement the comparison with the aforementioned two BNN-based BO approaches, BNN Laplace and BNN HMC, as specified in Section B.2. In addition, we also tested SBO with a single length-scale. We used the Matérn kernel. The results are reported in Fig. 6. Across all the benchmarks, the performance of the BNN based methods is worse than the median of GP based methods. In cases, such as SVM, Hartmann6 (300, 6) and Rosenbrock, the performance of the BNN based methods is close to or overlapping with the performance of the worst GP baselines. This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
Interestingly, even using a single length-scale parameter, the standard BO, namely SBO-Single — as long as it does not encounter gradient vanishing — still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-Single struggled in finding right locations to query, leading to poor performance.
C.2 Alternative Ways to GP Training
Figure 9: Final optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
Figure 10: Runtime optimization performance using a uniform prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
C.2.1 Maximum Likelihood Estimation (MLE)
In the main paper, we report the SBO performance when training GP via MAP estimation. We then examined another popular GP training approach, MLE training. We used ADAM optimization [17] [^2] We also experimented with using L-BFGS for MLE training. However, L-BFGS frequently encounters numerical issues and often fails to reliably complete model training. This is consistent with findings reported in [13]. The issue likely arises because, without constraints or regularization on parameters like the length-scales and amplitude, L-BFGS tends to explore too aggressively, leading to numerical instability. to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$. We employed ARD Matérn kernel. The ultimate and running optimization performance is shown in Fig.  9 and 10. As we can see, the BO performance of using MAP and MLE training is close, and in many cases significantly overlaps, such as in DNA, SVM, Ackley (150, 150) and StybTang (200, 200).
C.2.2 MAP Training with Alternative Priors
Second, we examined MAP training with alternative prior choices over the length-scale parameters. In the main paper, we use a uniform prior, Uniform(0.001, 30). In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular GP/BO libraries GPyTorch and BOTorch. To see how it is concentrated, in Fig. 11 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
We observe that the prior distribution Gamma(3, 6) places a large probability mass around its mean 0.5, strongly regularizing the length-scale parameters toward this value. According to our theoretical analysis (see Section 4), this introduces a significant risk of gradient vanishing in high-dimensional settings, potentially leading to failure in GP training.
To verify the risk, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180). As depicted in Fig. 12 (third column), the concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function. This observation aligns with findings reported in the SaasBO paper [6] (see Fig.1 of that paper).
We further investigated the impact of this prior within the optimization loop. As shown in Fig. 9, after placing the prior Gamma(3,6) over each length-scale parameter, the BO performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training.
Figure 11: Probability density curves. Gamma(2, 0.15) and Gamma(1.1, 0.5) are the priors for the amplitude and noise variance, respectively.
Figure 12: Alignment between the prediction (y-axis) and ground-truth function values (x-axis).
C.2.3 Fully Bayesian Training
Third, we investigated whether a fully Bayesian training approach would yield substantial improvements in BO. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP with ARD Matérn. We adopted the same settings as used in SaasBO-NUTS [6]: where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$. We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$. The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 7 , and Fig. 8. We also show the average running time per optimization step in Table 4 and Table 5. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal.
However, running NUTS is much more computationally costly compared to doing MAP and MLE. For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step. Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
C.3 Different Choice of $\lambda$ in UCB
Figure 13: Runtime optimization performance with different exploration level $\lambda$ in UCB.
We examined how the exploration parameter $\lambda$ in UCB influences the BO performance in high-dimensional optimization. To this end, we used the Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}. We conducted MLE training via ADAM optimization. We show the ultimate and runtime optimization performance in Fig. 13, respectively. It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close. However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$. This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
C.4 Alternative Acquisition Functions
Figure 14: Final performance with different acquisition functions.
Figure 15: Runtime performance with different acquisition functions.
Finally, we evaluated three other acquisition functions for high-dimensional optimization: Expected Improvement (EI), log-EI and Thompson sampling (TS). At each step, EI computes the expected improvement upon the best sample so far,
$\text{EI}({\bf x})=(\mu({\bf x})-f({\bf x}^{*}))\Psi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right)+\sqrt{v({\bf x})}\Phi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right),$ (16)
where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data. We used BOTorch to maximize EI through L-BFGS. The most recent work [1] introduced log-EI to overcome the numerical challenges in optimizing the original EI acquisition function. Accordingly, we also tested log-EI using the BoTorch implementation. To conduct TS at each step, we used Sobol Sequence [28] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior. We then selected the input with the largest function value as the next query. We denote this method by TS-3K. To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence. We denote this method as TS-10K. We used BOTorch implementation of both methods. We used the Matérn kernel and MLE training at each step.
The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig. 14 and 15, respectively. As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline. On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst. The failure of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1]. The failure of the TS might be due to the high-dimension of the input space. A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function. The use of log-EI performed excellently, achieving results comparable to those obtained with UCB.
Appendix D Limitation
Our current work has identified gradient vanishing as a critical failure mode of SBO in high-dimensional settings. We have proposed an effective, robust initialization method to mitigate this issue. However, we have not yet analyzed whether there is an upper limit to the input dimension for which SBO can function properly, or what that limit might be. To answer these questions, we will continue our investigation in future work, both theoretically and empirically.